# Relativistic Leaky Surface Plasmon Accelerator in Nanostructured Carbon Nanotube Forest

### v1

@@ II Relativistic surface plasmon and leaky field in the nanostructured CNT forest
We consider a $5\leavevmode\nobreak\ $\mathrm{T}\mathrm{W}$$ optical laser pulse of wavelength $\lambda_{L}=0.8\leavevmode\nobreak\ $\mu\mathrm{m}$$ passing through the vacuum channel of a nanostructured CNT forest target arranged in a porous pattern as seen in Fig. 1. The target is designed with a cylindrical vacuum channel of radius $r_{t}=3.0\leavevmode\nobreak\ $\mu\mathrm{m}$$. Details of the target configuration are summarised in Table.2 and discussed in Sec.IV.4. The laser pulse is linearly polarised in the $y$ direction and propagates in the $z$ direction. It is focused at the entrance of the channel $z=0\leavevmode\nobreak\ $\mu\mathrm{m}$$ with root-mean-square (RMS) waist $w_{0}=3.0\leavevmode\nobreak\ $\mu\mathrm{m}$$ and duration $\sigma_{\tau}=6.0\leavevmode\nobreak\ $\mathrm{f}\mathrm{s}$$. These parameters give a normalised laser strength $a_{0}=4.0$ or intensity $I=3.4\times 10^{19}\leavevmode\nobreak\ $\mathrm{W}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$$, with the field intensity of $1.46\times 10^{19}\leavevmode\nobreak\ $\mathrm{W}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$$ on the inner surface of the vacuum channel which drives the relativistic electron dynamics. The fully 3D simulations were conducted using the particle-in-cell (PIC) code WarpX [51].
@@
The wall electrons oscillate collectively driven by the ponderomotive scattering and are confined to the surface due to the transverse ionic potential, as shown in Fig.2(a), where the laser pulse has propagated $60\leavevmode\nobreak\ $\mu\mathrm{m}$$ within the forest channel. A microscale monochromatic leaky field inside the vacuum channel is excited at the relativistic resonant frequency, $\omega_{R}=\omega_{p}/\sqrt{2\gamma_{e}}$ and $\gamma_{e}$ the Lorentz factor of wall electron. The wavelength obtained from PIC, $\lambda_{R,\text{PIC}}=5.0\leavevmode\nobreak\ $\mu\mathrm{m}$$, agrees well with the theoretical prediction with relativistic effect, $\lambda_{R,\text{Theo.}}=2\pi c/\omega_{R}=4.8\leavevmode\nobreak\ $\mu\mathrm{m}$$, as discussed in Sec. IV.2. In comparision, the non-relativistic prediction is $\lambda_{r,\text{Theo.}}=2\pi c/\omega_{r}=3.3\leavevmode\nobreak\ $\mu\mathrm{m}$$ [4]. The energy of RSP is $\hbar\omega_{R}=0.24\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$\gg E_{F}=0.05\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$, where $E_{F}=\hbar^{2}(3\pi^{2}n_{e})^{2/3}/2m_{e}$ is the Fermi energy. This indicates that Landau damping does not occur [52]. On the other side, as the RSP energy is slightly higher than that of optical phonon in typical CNTs in the range of $0.15-0.2\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$ [53], the optical damping effect can also be neglected. As a result, a long propagation distance of RSP is possible. For example sub-$\mathrm{m}\mathrm{m}$ distance has been seen from PIC simulations. The longitudinal component of the leaky field remains quasi-uniform across the vacuum channel with an amplitude of $0.2\leavevmode\nobreak\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$$, providing periodic accelerating phases for both negatively and positively charged particles, as shown in Fig. 2 (a), (c), and (d), which is advantageous for maintaining beam quality. The transverse component of the field, $W_{x}=E_{x}-cB_{y}$, supplies a focusing force for positrons, phase-matched with the corresponding acceleration phase, as shown in Fig.2(b) and by the black line in (e). This field reaches an amplitude of $0.3\leavevmode\nobreak\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$$ on the surface and decays transversely. In the phase for electron acceleration, the transverse field vanishes in the vacuum region, instead providing a confining force on the inner surface, as shown by the green line in Fig. 2(e). Consequently, electrons can also be confined within the vacuum channel and accelerated. The laser pulse propagates through the vacuum channel without any noticeable deformation, as shown in Fig. 2 (b) and (f). This demonstrates that the scheme is dynamically stable and can be effectively scaled for long porous targets, limited only by the depletion of the laser energy.
@@
In order to be successfully accelerated, a relativistic electron or positron beam should be injected into the correct phase. We now put either an electron or positron beam in the corresponding accelerating phase, indicated by vertical dashed pink and black lines in Fig. 2(a) and (b), respectively. Both of the beams have the same parameters except for the negative or positive charge, as summarised in Table.1.
@@
The evolution of energy gain and transverse beam size is shown in Fig. 3 (a) and (b), respectively. All the electrons and positrons inside the witness beam can be preserved and accelerated which indicates that the transverse component of the RSPs on the inner surface can sufficiently confine both beams. With the positron beam, after $z=150\leavevmode\nobreak\ $\mu\mathrm{m}$$ propagation, the beam gains energy $\Delta E=19\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$ energy which indicates the mean energy gain rate as $R_{E}=0.12\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$. The beam size is reduced by half due to the strong transversely leaky field and is then equilibrated at $z=130\leavevmode\nobreak\ $\mu\mathrm{m}$$, where the focusing field of RSP is in balance with the self space-charge field of the witness beam. For the witness electron beam, the energy gain is $\Delta E=25\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$, which indicates the mean energy gain rate as $R_{E}=0.16\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$. The radial beam size is preserved with a small increase as the transverse component of the leaky field at the acceleration phase for the electron beam is weaker than that for the positron beam as shown in Fig. 2(b) and (e). Additionally, the quasi-linear energy gain indicates that the phase slippage between the witness beam and leaky field is slow and then enables the extensible acceleration for high-power laser pulse as discussed in Sec. IV.3.
@@ III.1 Edge injection
Figure 4: PIC results: Evolution of the electron density (grey colourmap), longitudinal electric field (blue-red colourmap), and laser field (green-yellow colourmap) at five propagation positions: $z=18,21,24,27$ and $30\leavevmode\nobreak\ $\mu\mathrm{m}$$. The laser strength is $a_{0}=5.0$, with other parameters the same as those in Fig. 2. The horizontal green dashed lines indicate the positions of the inner surface of the forest.
With a low-intensity laser pulse, wall electrons gain relatively low momentum and are unable to overcome the two-dimensional (2D) transverse confinement on the channel surface, preventing them from freely crossing the inner vacuum surface. However, the situation changes during the laser enters the forest channel. Here, the vertical edge provides an additional dimension of confinement in the longitudinal direction, which compresses the streaming electrons, as shown in Fig. 4. This compression allows some electrons to gain additional momentum which allows them to overcome the transverse confinement and emit into the vacuum channel. Since the RSPs are localised waves, the emitted electrons are naturally in the acceleration phase along the surface and are simultaneously accelerated longitudinally while drifting transversely through the vacuum channel. These electrons collide at the channel centre and become trapped in the leaky field once they acquire sufficient longitudinal momentum. This method is analogous to catapult injection at the edge of stacked graphene layers [55]. PIC simulations reveal that after $100\leavevmode\nobreak\ $\mu\mathrm{m}$$ of propagation, a $42\leavevmode\nobreak\ $\mathrm{p}\mathrm{C}$$ electron beam with a duration of $3.3\leavevmode\nobreak\ $\mathrm{f}\mathrm{s}$$ is accelerated to $35\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$ with a $1\%$ energy spread, corresponding to a mean acceleration gradient of $0.35\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$. The beam transverse size is $\sigma_{r}=0.2\leavevmode\nobreak\ $\mu\mathrm{m}$$. Such an ultrashort, high-quality beam holds unprecedented potential for advanced applications in high-field science and beyond [56].
@@ III.2 Self injection
Figure 5: PIC results: (a) Electron density distribution ($n_{e}$, grey colourmap), and longitudinal electric field ($E_{z}$, blue-red colourmap). (b) Transverse field, $W_{x}$ (blue-red colourmap) and laser field (yellow-green colourmap). (c) Phase space of the accelerated electron beam after exiting the $80\leavevmode\nobreak\ $\mu\mathrm{m}$$ long forest target. (d) Energy spectrum corresponding to (c). All the laser and target parameters are the same as these used for Fig. 2 except the laser strength which is increased to $a_{0}=10$. The horizontal green dashed lines in (a) and (b) indicate the positions of the inner surface of the forest.
@@
With increased laser strength, electron oscillations on the surface intensify, leading to a significant enhancement of the RSP amplitude, as demonstrated in Fig.5(a), where a laser pulse with $a_{0}=10$ or peak power $P=30\leavevmode\nobreak\ $\mathrm{T}\mathrm{W}$$ is used. This enhancement enables some wall electrons to gain sufficient transverse momentum to cross the inner surface and penetrate deeply into the vacuum region. These electrons naturally traverse the forest surface through the accelerating phase of the RSP and align with the accelerating phase of the leaky field in the vacuum channel. The leaky field, with a magnitude as high as $0.6\leavevmode\nobreak\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$$, facilitates the rapid longitudinal acceleration of these transiting electrons, allowing them to catch up with the leaky field within just a few $\mu\mathrm{m}$. The transverse field of the RSP acts as a confining force along the inner surface, preventing the transited electrons from escaping back into the forest wall, as illustrated in Fig. 5(b). Here, the transverse component of the leaky field vanishes across most of the acceleration phase region, ensuring stability. As a result, a substantial number of electrons are trapped and continue to be efficiently accelerated.
@@
A $204\leavevmode\nobreak\ $\mathrm{p}\mathrm{C}$$ electron beam is self-injected and accelerated to $35\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$ over an $80\leavevmode\nobreak\ $\mu\mathrm{m}$$ long-distance, corresponding to a peak acceleration gradient of $G=0.43\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$. The mean beam energy is $20.4\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$, yielding a mean acceleration gradient of $G=0.26\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$. The phase space and energy spectrum of the accelerated beam are presented in Fig.5(c) and (d), respectively, which show a broad spectrum. Electrons continuously transit across the inner surface until the laser pulse is sufficiently depleted, allowing the self-injection process to persist, as discussed in Sec.IV.3. The high-energy component peaks at $30\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$ with a $26\%$ spread, while the low-energy component peaks at $5.4\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$ with a $1\%$ spread. This self-injection mechanism is fully feasible with currently available high-power laser facilities, offering a promising pathway for experimental demonstration of this proposed CNT forest-based plasma accelerator.
IV Discussion and methods
IV.1 Quantum degeneracy and correlation
The excitation of RSPs discussed in this paper relies on the collective dynamics of high-density conduction electrons which can be treated as a classical plasma if the CNT forest is in quantum degenerate state [1]. To establish the feasibility of this condition, we first examine how it can be satisfied under typical experimental conditions, ensuring the required electron density and temperature ranges are achievable within current experimental setups.
Metallic crystals such as CNTs exhibit quantum degeneracy at low temperatures or high carrier densities, a state governed by the Pauli exclusion principle [57, 58]. In this quantum degenerate state, the conduction electrons act as highly degenerate carriers, forming a Fermi gas that dictates the quantum behaviour of the fermions, such as minimum scattering. Electrons with energy near the Fermi level can move freely in response to small perturbations, such as electric fields, resulting in high electrical conductivity and low thermal sensitivity. The quantum degeneracy parameter, $\theta$, quantifies the extent to which a system is quantum degenerate and is defined by
$\theta=\frac{E_{F}}{K_{B}T}\mathrm{.}$ (1)
$K_{B}$ is the Boltzmann constant and $T$ is the temperature. Quantum degeneracy requires $\theta\gg 1$, indicating that the Fermi energy is significantly larger than the thermal energy. In CNTs, the conduction electron density ranges from $10^{18}-10^{24}\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}^{-3}$$ and can support a degenerate Fermi electron gas over a wide parameter space (shown in Fig. 6). For example, with an electron density of $n_{e}=2.0\times 10^{20}\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}^{-3}$$, quantum degeneracy is maintained at temperatures up to $T=400\leavevmode\nobreak\ $\mathrm{K}$$. Therefore, under typical laboratory conditions, these conduction electrons form a collisionless dense plasma environment in which the positive background charge can be assumed to be uniformly distributed, creating a stabilising field. Furthermore, interactions between electrons and lattice vibrations and photons are often neglected. This validates the classical kinetic method for studying the RSP excitation in the CNT forest, as presented in this paper.
Figure 6: Quantum degeneracy parameter $\theta$ as a function of temperature $T$ and electron density $n_{e}$.
In an electron gas, particles interact strongly due to the long-range nature of the Coulomb force. When the electron density is sufficiently high, collective behaviour, such as plasmons, becomes dominant. In this regime, the dynamics of each electron do not depend individually on the other electrons but on a smeared potential created by the collective electron cloud. This correlated state of the electron gas leads to macroscopic quantum effects and collective behaviour [1]. To sustain collective behaviour, the damping time must exceed the average oscillation period of these collective modes, which is given by the ratio
$\eta=\frac{\tau_{\text{damp}}}{\tau_{\text{coll}}}=\frac{\hbar/(2\pi KK_{B}T)}{2\pi/\omega_{p}}\mathrm{,}$ (2)
where $\tau_{\text{damp}}$ is the damping time of the degenerate electron gas and $\tau_{\text{coll}}$ is the period of the collective oscillations. Here $K$ represents the Luttinger parameter, typically in the range of $0.2$ to $0.5$ for metallic CNTs [59]. Eq. (2) indicates that quantum correlations are also favoured under conditions of high electron density or low temperature. In the case of CNTs with typical parameters, the high value of $\eta\gg 1$ confirms the presence of significant quantum correlations in this system.
As a result, the conduction electron gas in CNTs can generally be approximated as a classical plasma in which electrons move freely in a quasi-uniform background. The dominant collective dynamics in this regime can be effectively studied using classical kinetic theory. Quantum effects such as plasma dispersion corrections, photon-electron coupling, polariton interactions and collision damping can be safely neglected due to their minimal influence under these conditions.
IV.2 Leaky field of RSPs in flat and cylindrical geometry
Figure 7: PIC results: SPs excitation on the flat CNT surface. (a) Density distribution of wall electrons (grey colourmap) and laser field (green-yellow colourmap). (b) Longitudinal acceleration field $E_{z}$. (c) Line plots of $E_{z}$ along $x=0\leavevmode\nobreak\ $\mu\mathrm{m}$$ (black) and $x=-2.5\leavevmode\nobreak\ $\mu\mathrm{m}$$ (green), indicated by the horizontal dashed lines in (b), respectively. (d) Line plots of $E_{z}$ along $z=46\leavevmode\nobreak\ $\mu\mathrm{m}$$ (orange), $z=48\leavevmode\nobreak\ $\mu\mathrm{m}$$ (green) and $z=50\leavevmode\nobreak\ $\mu\mathrm{m}$$ (black), indicated by the vertical dashed lines in (b) respectively. The vertical grey line indicates the position of the interface between the CNT wall and the vacuum. The corresponding theoretical estimations are shown by the dashed lines, respectively.
A high-intensity laser pulse propagating along a plasma-vacuum boundary excites the confined surface waves which can be treated as the polarisation of the plasma in response to the laser field [49]. The dispersion of RSP with relativistic effects is given by
$1+\varepsilon\sqrt{\frac{n^{2}-1}{n^{2}-\varepsilon}}=0\mathrm{,}$ (3)
where $\varepsilon=1-\omega_{p}^{2}/\gamma_{e}\omega_{sp}^{2}$ and $\gamma_{e}$ is the Lorentz factor of the electron. $n=ck_{z}/\omega_{sp}=c/v_{\text{ph}}$ with $v_{\text{ph}}$ the phase velocity of the RSP and $k_{z}$ the longitudinal component of RSP wavevector. $\varepsilon=-1$ in resonance excitation [6]. Eq. (3) can only be satisfied if $n>1$. $n>1$ indicates that the phase velocity of the RSP is less than $c$, leading to the dephasing problem. The phase velocity of the RSP with resonance condition is given as
where $\alpha=2\gamma_{e}-1$. In the non-relativistic case $\gamma_{e}\simeq 1$, $v_{\text{ph}}\simeq 0$, indicating that SP is a standing wave with a dominant electrostatic component [49] and is unsuitable for relativistic particle acceleration. In the ultrarelativistic case $\gamma_{e}\gg 1$, $v_{\text{ph}}$ approaches $c$, which allows the acceleration of relativistic particles. However, when the laser strength is high enough, the wall electrons are no longer confined to the surface and can move freely to form the bubble wakefield, similar to that in a gaseous plasma [40]. The advantage of this type of bubble is that it is dephasingless since the group velocity of the laser pulse is equal to the speed of light in vacuum $c$.
For a Gaussian pulse propagating along the axis, the maximum $\gamma_{e}$ is achieved at $r_{L}=r_{t}$ and then $a(r_{t})=a_{0}e^{-r_{t}^{2}/w_{0}^{2}}$, which suggests that the ratio $r_{t}/w_{0}$ also plays a critical role in RSP excitation. The relativistic effect requires the channel radius $r_{t}<w_{0}\sqrt{\ln(a_{0})}$. With nanostructured CNT forests of flexible density, the RSP can be directly excited independent of laser polarisation or the matching medium. At resonance, the RSP exhibits a dominant localised electrostatic component, with its phase velocity closely matching the group velocity of the laser pulse, enabling efficient energy transfer and interaction.
We first consider a CNT forest target with a flat surface where RSP can be excited by paraxially passing a high-intensity laser pulse along $z$-direction, as shown in Fig.7(a). Here, the surface is positioned at $x=-3.0\leavevmode\nobreak\ $\mu\mathrm{m}$$ and extends infinitely in $y$ and $z$ directions, with a wall density of $n_{e}=2\times 10^{20}\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}^{-3}$$. The EM field with an amplitude of $0.1\leavevmode\nobreak\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$$ can leak into a small region of the surface and decay exponentially, as shown in Fig. 7(b) to (d). Theoretically, this behaviour can be explained by solving Maxwell’s equations with appropriate boundary conditions at the surface [5]. The longitudinal electric field can be expressed as
$E_{z}=E_{z0}e^{-k_{x}x}e^{i(k_{z}z-\omega_{sp}t)}\mathrm{,}$ (5)
where $E_{z0}$ is the amplitude and $k_{x}=\omega_{sp}\sqrt{n^{2}-1}/c$. The resonance wavelength is theoretically calculated to be $\lambda_{R}=3.24\leavevmode\nobreak\ $\mu\mathrm{m}$$, which agrees well with the simulation result $\lambda_{sp,PIC}\simeq 3.3\leavevmode\nobreak\ $\mu\mathrm{m}$$, as shown in Fig. 7(c). RSP decays slower in vacuum than in CNTs, since $k_{x,\text{vacuum}}=\omega_{sp}\sqrt{n^{2}-1}/c<k_{x,\text{CNTs}}=\omega_{sp}\sqrt{n^{2}-\varepsilon}/c$ [49]. The longitudinal component $E_{z}$ of the RSP along the $x$-direction is plotted at three different positions by solid lines in Fig.7(d) and shows good agreement with the theoretical estimates indicated by the dashed lines. The observed deviations, e.g. at $z=50\leavevmode\nobreak\ $\mu\mathrm{m}$$ as shown by the solid and dashed black lines, are due to the influence of the laser field. With the evanescent nature of the leaky field on the flat surface, the accelerating volume is limited to a small region, e.g. $|\Delta x|<1/k_{sp}\leavevmode\nobreak\ \sim 0.2\leavevmode\nobreak\ $\mu\mathrm{m}$$ for our case, making the acceleration highly sensitive to the angular distribution of both the driving and witness beams. Consequently, the total charge and energy gain of the witness beam are limited.
The CNT forest can be nanostructured to include a $\mu\mathrm{m}$-scale cylindrical vacuum channel by arranging the CNTs in an array pattern. This configuration theoretically breaks the cylindrical symmetry due to inter-tube Coulomb coupling among the periodically arranged CNTs [60], which can result in high-frequency leaky fields within the vacuum channel. However, the porous targets consist of hundreds of CNTs, with bulk dimensions significantly larger than those of a single CNT. In this case, the inter-structural interactions can be approximated as effectively uniform [40]. Consequently, mode coupling is not significant, as illustrated in Fig. 2. Without couplings between SP modes of different azimuthal indices, the leaky field can remain monochromatic, as seen in Fig. 2(a) and (b). When the intensity of the laser field is sufficiently high, these couplings become significant, potentially introducing high-frequency components into the leaky field, as shown in Fig. 5(a) and (b). Therefore, there is a compromise between the field strength and the field distribution. The quality of the leaky field, including field structure and the volume available for acceleration, can be significantly enhanced by employing a radially symmetric target, such as a hollow nanostructured CNT forest, as depicted in Fig. 1.
In a nanostructured CNT channel with a large radius $r_{t}$ (e.g., $k_{sp}r_{t}\gg 1$), the eigenfrequencies of all SP modes become less sensitive to surface curvature and approach the resonant frequency [61, 62]. Since the length of the target is significantly larger than the channel radius ($L\gg r_{t}$), the electric potential of leaky RSP modes in a cylindrical channel can be derived by solving the Laplace equation in cylindrical coordinates $(r,\phi,z)$ as follows [37]:
which decays to zero and exists only in a narrow region near the surface, in agreement with the simulation results shown in Fig. 2(b) and 5(b). From Eqs. (7) and (8), it can be seen that there is a $\pi/2$ phase region where either negatively or positively charged particles can experience both acceleration and focusing forces. Therefore, the leaky field of the fundamental RSP mode, $m=0$, is suitable for accelerating both electrons and positrons.
IV.3 Wall electron transition and cascaded leaky RSPs excitation

### v2

@@ II.2 On cylindrical surface
The structure of the leaky EM field structure becomes different on a cylindrical surface which can be nanostructured into the CNT forest as shown in Fig. 1, where the radius of the surface is $r_{t}$. With bulk dimensions significantly larger than those of a single CNT, the inter-structural interactions between single CNT can be approximated as effectively uniform [31]. Without couplings between SP modes of different azimuthal indices, the leaky field can remain monochromatic with the resonant frequency, as seen in Fig. 3(a) and (b). The PIC simulations are done with a $5\leavevmode\nobreak\ $\mathrm{T}\mathrm{W}$$ Gaussian laser pulse same as that used in Fig. 2 passing through the vacuum channel of a nanostructured CNT forest target arranged in a porous pattern as seen in Fig. 1. The target is designed with a cylindrical vacuum channel of radius $r_{t}=3.0\leavevmode\nobreak\ $\mu\mathrm{m}$$. Details of the target configuration are summarised in Table.1. The numerical parameters are also the same as those in Fig. 2.
Table 1: Parameters of the porous CNT target used in PIC simulations.
Layer gap, $g$ $50\leavevmode\nobreak\ $\mathrm{n}\mathrm{m}$$
Pore diameter, $d$ $100\leavevmode\nobreak\ $\mathrm{n}\mathrm{m}$$
Forest channel radius, $r_{t}$ $3\leavevmode\nobreak\ $\mu\mathrm{m}$$
Filling factor, $f_{f}$ $0.65$
Number of layers, $N$ $20$
@@
SP excitation and the leaky field in the vacuum channel are shown in Fig. 3. The wall electrons oscillate collectively driven by the ponderomotive scattering with the laser pulse and are confined to the surface due to the transverse ionic potential, as shown in Fig. 3(a), where the laser pulse has propagated $60\leavevmode\nobreak\ $\mu\mathrm{m}$$ within the forest channel. A microscale monochromatic leaky field inside the vacuum channel is excited. The wavelength obtained from PIC, $\lambda_{sp,\text{PIC}}=5.0\leavevmode\nobreak\ $\mu\mathrm{m}$$, agrees well with the theoretical prediction with relativistic effect, $\lambda_{sp,\text{Theo.}}=2\pi c/\omega_{sp}=4.8\leavevmode\nobreak\ $\mu\mathrm{m}$$. The longitudinal component of the leaky field remains quasi-uniform across the vacuum channel with an amplitude of $0.2\leavevmode\nobreak\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$$, providing periodic accelerating phases for both negatively and positively charged particles, as shown in Fig. 3 (a), (c), and (d), which is advantageous for maintaining beam quality. The transverse component of the field, $W_{x}=E_{x}-cB_{y}$, supplies a focusing force for positrons, phase-matched with the corresponding acceleration phase, as shown in Fig. 3(b) and by the black line in (e). This field reaches an amplitude of $0.3\leavevmode\nobreak\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$$ on the surface and decays transversely. In the phase for electron acceleration, the transverse field vanishes in the vacuum region, instead providing a confining force on the inner surface, as shown by the green line in Fig. 3(e). Consequently, electrons can also be confined within the vacuum channel and accelerated. The laser pulse propagates through the vacuum channel without any noticeable deformation, as shown in Fig. 3(b) and (f). This demonstrates that the scheme is dynamically stable and can be effectively scaled for longer targets, limited only by the depletion of the laser energy.
Theoretically, we can understand harmonic wakefield dynamics with the assumption that a channel radius $r_{t}$ is large, as $k_{sp}r_{t}\gg 1$. In this case, the eigenfrequencies of all SP modes become less sensitive to surface curvature and approach the resonant frequency [32, 33]. Since the length of the target is significantly larger than the channel radius ($L\gg r_{t}$), the electric potential of leaky SP modes in a cylindrical vacuum channel can be derived by solving the Laplace equation in cylindrical coordinates $(r,\phi,z)$ as follows [34]:
which decays to zero and exists only in a narrow region near the surface, in agreement with the simulation results shown in Fig. 3(b) and 6(b). From Eqs. (5) and (6), it can be seen that there is a $\pi/2$ phase region where either negatively or positively charged particles can experience both acceleration and focusing forces. Therefore, the leaky field of the fundamental mode, $m=0$, is suitable for accelerating both electrons and positrons.
@@
To be successfully accelerated, a relativistic electron or positron beam should be injected into the correct phase. We now put either an electron or positron beam in the corresponding accelerating phase, indicated by vertical dashed pink and black lines in Fig. 3(a) and (b), respectively. Both of the beams have the same parameters except for the negative or positive charge, as summarised in Table 2.
@@
The evolution of energy gain and transverse beam size is shown in Fig. 4 (a) and (b), respectively. All the electrons and positrons inside the witness beam can be preserved and accelerated which indicates that the transverse component of the leaky field can sufficiently confine both beams. With the positron beam, after $z=150\leavevmode\nobreak\ $\mu\mathrm{m}$$ propagation, the beam gains energy $\Delta E=19\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$ energy which indicates the mean energy gain rate as $R_{E}\approx 0.13\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$. The beam size is reduced by half until being equilibrated at $z=130\leavevmode\nobreak\ $\mu\mathrm{m}$$, where the focusing field is in balance with the self-space-charge field of the witness beam. For the witness electron beam, the energy gain is $\Delta E=25\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$, which indicates the mean energy gain rate as $R_{E}\approx 0.17\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$. The radial beam size is preserved with a small increase as the transverse component of the leaky field at the acceleration phase for the electron beam is weaker than that for the positron beam as shown in Fig. 3(b) and (e). Additionally, the quasi-linear energy gain indicates that the phase slippage between the witness beam and leaky field is slow and then enables the extensible acceleration for high-power laser pulse.
@@ III.1 Edge injection
Figure 5: PIC results: Evolution of the electron density (grey colourmap), longitudinal electric field (blue-red colourmap), and laser field (green-yellow colourmap) at five propagation positions: $z=18,21,24,27$ and $30\leavevmode\nobreak\ $\mu\mathrm{m}$$. The laser strength is $a_{0}=5.0$, with other parameters the same as those in Fig. 3. The horizontal green dashed lines indicate the initial positions of the inner surface of the forest.
With a low-intensity laser pulse, wall electrons gain relatively low momentum and are unable to overcome the two-dimensional (2D) transverse confinement on the channel surface, preventing them from freely crossing the inner vacuum surface. However, the situation changes during the laser enters the forest channel. Here, the vertical edge provides an additional dimension of confinement in the longitudinal direction, which compresses the streaming electrons, as shown in Fig. 5. This compression allows some electrons to gain additional momentum which allows them to overcome the transverse confinement and emit into the vacuum channel. Since the SPs are localised waves, the emitted electrons are naturally in the acceleration phase along the surface and are simultaneously accelerated longitudinally while drifting transversely through the vacuum channel. These electrons collide at the channel centre and become trapped in the leaky field once they acquire sufficient longitudinal momentum. This method is analogous to catapult injection at the edge of stacked graphene layers [36]. PIC simulations reveal that after $100\leavevmode\nobreak\ $\mu\mathrm{m}$$ of propagation, a $42\leavevmode\nobreak\ $\mathrm{p}\mathrm{C}$$ electron beam with a duration of $3.3\leavevmode\nobreak\ $\mathrm{f}\mathrm{s}$$ is accelerated to $35\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$ with a $1\%$ energy spread, corresponding to a mean acceleration gradient of $0.35\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$. The beam transverse size is $\sigma_{r}=0.2\leavevmode\nobreak\ $\mu\mathrm{m}$$. Such an ultrashort, high-quality beam holds unprecedented potential for advanced applications in high-field science and beyond [37].
@@ III.2 Self injection
Figure 6: PIC results: (a) Electron density distribution ($n_{e}$, grey colourmap), and longitudinal electric field ($E_{z}$, blue-red colourmap). (b) Transverse field, $W_{x}$ (blue-red colourmap) and laser field (yellow-green colourmap). (c) Phase space of the accelerated electron beam after exiting the $80\leavevmode\nobreak\ $\mu\mathrm{m}$$ long forest target. (d) Energy spectrum corresponding to (c). All the laser and target parameters are the same as those used for Fig. 3 except the laser strength which is increased to $a_{0}=10$. The horizontal green dashed lines in (a) and (b) indicate the initial positions of the inner surface of the forest.
@@
With increased laser strength, electron oscillations on the surface intensify, leading to a significant enhancement of the SP amplitude, as demonstrated in Fig. 6(a), where a laser pulse with $a_{0}=10$ or peak power $P=30\leavevmode\nobreak\ $\mathrm{T}\mathrm{W}$$ is used. This enhancement enables some wall electrons to gain sufficient transverse momentum to cross the inner surface and penetrate deeply into the vacuum region. These electrons naturally traverse the forest surface through the accelerating phase of the SP and align with the accelerating phase of the leaky field in the vacuum channel. The leaky field, with a magnitude as high as $0.6\leavevmode\nobreak\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$$, facilitates the rapid longitudinal acceleration of these transiting electrons, allowing them to catch up with the leaky field within just a few $\mu\mathrm{m}$. The transverse field of the SP acts as a confining force along the inner surface, preventing the transited electrons from escaping back into the forest wall, as illustrated in Fig. 6(b). As a result, a substantial number of electrons are trapped and continue to be efficiently accelerated.
@@
A $204\leavevmode\nobreak\ $\mathrm{p}\mathrm{C}$$ electron beam is self-injected and accelerated to $35\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$ over an $80\leavevmode\nobreak\ $\mu\mathrm{m}$$ long-distance, corresponding to a peak acceleration gradient of $G\approx 0.44\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$. The mean beam energy is $20.4\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$, yielding a mean acceleration gradient of $G\approx 0.26\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$. The phase space and energy spectrum of the accelerated beam are presented in Fig. 6(c) and (d), respectively, which show a broad spectrum. Electrons continuously transit across the inner surface until the laser pulse is sufficiently depleted, allowing the self-injection process to persist as discussed later. The high-energy component peaks at $30\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$ with a $26\%$ spread, while the low-energy component peaks at $5.4\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$ with a $1\%$ spread. This self-injection mechanism is fully feasible with currently available high-power laser facilities, offering a promising pathway for experimental demonstration of this proposed CNT forest-based plasma accelerator.